A bottle rocket is launched from the ground with a velocity of 30.0 m/s upward.
a) What is the maximum height of the bottle rocket? ________________________
b) How long will it take the bottle rocket to reach its max height? ________________________
c) How long will it take the bottle rocket to land? ________________________
d) What will be the bottle rocket’s velocity at landing?
1
Expert's answer
2020-06-01T14:24:04-0400

Given:

"{v_0=30.0\:\rm m/s;} \\
g=9.81\;\rm m/s^2."

a)


"h_{\rm max}=\frac{v_0^2}{2g}=\frac{30.0^2}{2\times 9.81}=45.9\:\rm m"

b)

"t_1=\frac{v_0}{g}=\frac{30.0}{9.81}=3.06\:\rm s"

c)

"t_2=2t_1=2\times 3.06=6.12\:\rm s"

d)

"v_f=v_0=30.0\:\rm m/s \quad (downward)"